OpenStudy (anonymous):

Milton Jones scored 88 and 93 on his first two tests in Buisness Management class. What score must he get on his third test to keep at least an average of 84 or greater?

OpenStudy (anonymous):

\[\frac{83+93+x}{3}\geq 84\] is a start

OpenStudy (anonymous):

or maybe just \[\frac{176+x}{3}\geq 84\]

OpenStudy (anonymous):

then multiply both sides by 3, and finally subtract \(176\) so get \(x\)

OpenStudy (anonymous):

this is a lot of different types of questions. must be a strange class

OpenStudy (anonymous):

for that question, i got 76

OpenStudy (anonymous):

yeah this class is a pain in my retrice.. i'm HORRIBLE at math

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

but these are my only options .. 90 or higher 84 or higher 71 or higher 77 or higher

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

ok i think i made another typo, try \[\frac{88+94+x}{3}\geq 84\] \[\frac{181+x}{3}\geq 84\] \[181+x\geq 252\] \[x\geq 71\]

OpenStudy (anonymous):

i guess you are not that horrible in math, i am not so good in reading

OpenStudy (anonymous):

go with "71 or higher" as usual it is option C

OpenStudy (anonymous):

yeah you are right. i guess you are not that horrible in math